I need to use the discriminant to determine the number and type of solutions for this equation: x^2+2x=-11
Thanks!

Answer by stanbon(75887)

I need to use the discriminant to determine the number and type of solutions for this equation: x^2+2x=-11
-----
a = 1
b = 2
c = 11
-----
b^2-4ac = 2^2-4*1*11 = 4-44 < 0
---
Since the discriminant is negative there are 2 complex number solutions.
------------
Cheers,
Stan H.
==============